Grade 11Organic Chemistry

A sugar syrup of weight 214.2g contains 34.2g of sugar . Calculate the molar concentration and mole fraction of sugar in syrup

Sundip
8 Years agoGrade 11

2 Answers

Adarsh
8 Years ago
Dear,
342g of sugar contain 1 mole of sugar so 34.2g contain 0.1 mole of sugar ..
weight of water in solution = 180gm = 10 mole of water 
mole fraction of sugar =  10+0.1/10= 1.01
Thanks and Regards 
Adarsh
RIYA
8 Years ago
Weight of sugar syrup = 214.2g
Weight of sugar present= 34.2
Molecular mass of sugar(C12H22O11)=342
(I)Calculation of molality:
Number of moles of sugar=34.2/342=0.1
Weight of water =214.2-34.2=180g=0.18kg
So; molality of the solution = No. Of moles of sugar/Weight of water (in kg)
0.1/0.18=0.56mol/kg
Hence; the molality concentration of the given soln is 0.56m.
(ii) calculation of moles fraction :
Number of moles of sugar (n1)=0.1
Number of moles of water (n2)=180/18=10
Total number of moles in solution =n1+n2=0.1+10=10.1
So; The mole fraction of sugar n1/n1+n2=0.1/10.1=9.9/1000.